### v2

# Operad of Posets 101: The Wixárika Posets

Abstract.
We study sets whose combinatorics are related to the combinatorics of posets.
The language of operads provides us with tools to better understand the combinatorics of these objects. In this note we describe a non-trivial example of a suboperad, called the Wixárika posets, together with its associated algebras. This example is rich enough to showcase the particularities of the field, without delving into technicalities.
1. Introduction
This note is intended as an extension of lecture notes designed to assist students in understanding the theory of operads, with a particular focus on algebras over the operad of posets. This pedagogical aim is inspired by the well-known role of Abdus Salam as both a teacher and a promoter of science, emphasizing the importance of making advanced mathematical structures accessible to students and early-career researchers. Accordingly, this note does not contain new results.
Let us begin by providing an overview of what this note is about. Let $P$ be a partially ordered set (poset). We associate to $P$ a directed acyclic simple graph, called the Hasse diagram $H_{P}$ of $P$: Each vertex of $H_{P}$ represents an element of $P$ and there is an arrow from $x$ to $y$ if $x<y$ and there is no $z$ with $x<z<y$. We can visualize compositions of posets $P$ and $Q$ in terms of their Hasse diagrams. Given a point $p\in P$, the composition $P\circ_{p}Q$ is the poset whose Hasse diagram looks like the Hasse diagram of $P$ except that the vertex $p$ is replaced by the diagram $H_{Q}$. Using this composition we can associate to the poset $\{x,y\}$, the disjoint union of two elements, a map of posets that sends $P,Q$ to $\{x,y\}(P,Q)=(\{x,y\}\circ_{x}P)\circ_{y}Q=P\sqcup Q$, the disjoint union of $P$ and $Q$.
To formalize the algebraic structure highlighted in the previous example we use the language of operads. An operad consists of sets $O(n)$ for all $n\geq 1$, and composition maps $O(n)\times O(k_{1})\times\cdots\times O(k_{n})\mapsto O(\sum_{i=1}^{n}k_{i})$, satisfying certain compatibility conditions (see Section 4). In the previous example, the operad of finite posets is formed by $\mathrm{Poset}(n)$, the sets of posets with $n$ elements. For $P\in\mathrm{Poset}(n),Q_{i}\in\mathrm{Poset}(k_{i}),i=1,\cdots,n,$ the composition $P\circ(Q_{1},\cdots,Q_{n})$ is the poset whose Hasse diagram is obtained by replacing for each $i=1,\ldots,n$ the $i$-th vertex of $H_{P}$ by $H_{Q_{i}}$.
Motivated by this construction we aim to study sets $X$ together with maps $\mathrm{Poset}(n)\mapsto\mathrm{Map}(X^{n},X)$. For a correspondence that sends a poset $P$ to a map $f_{P}$, we also require that the poset $P\circ_{p_{i}}Q$ is send to $f_{P}(id,\cdots,id,f_{Q},id,\cdots,id)$ where $f_{Q}$ is located on the $i$-th entry.
Here is an explicit example: R. Stanley [STA70] defined the strict $n$-order polynomials $\Omega^{\circ}(P,n)$ of $P$ as the number of strict order-preserving maps from $P$ to the ordered set $\langle n\rangle=\{1<\cdots<n\}$. By definition $\Omega^{\circ}(P\sqcup Q,n)=\Omega^{\circ}(P,n)\Omega^{\circ}(Q,n)$. We interpret the previous result as suggesting that the endomorphism associated with $\{x,y\}$ is the product of order polynomials, that is, $\{x,y\}(\Omega^{\circ}(P,n),\Omega^{\circ}(Q,n))=\Omega^{\circ}\left(\{x,y\}(P,Q),n\right)=\Omega^{\circ}(P,n)\Omega^{\circ}(Q,n)$. By abuse of notation we denote with $\{x,y\}$ an operation on polynomials and an operation on posets, since this choice clarifies the algebra structure.
Another example of a set in which posets act as endomorphisms comes from polytope theory. The order polytope of a poset $P$ is defined as
$\mathrm{Poly}(P)=\{f\colon P\longrightarrow[0,1]\mid x\leq_{P}y\Rightarrow f(x)\leq f(y)\}.$
One can verify that $\mathrm{Poly}(\{x,y\}(P,Q))=\{x,y\}(\mathrm{Poly}(P),\mathrm{Poly}(Q))$ if $\{x,y\}$ evaluated on polytopes is the Minkowski sum of polytopes.
From the operadic point of view, we study sets in which the elements of the operad define maps. These sets are called algebras over the operad. Above we described order polynomials and order polytopes as examples of algebras over the operad of posets. It will become clear that posets themselves are an algebra over the operad of posets, and we will learn how to transfer information from one algebra to the other.
As this is an introductory text, it will be more convenient to work with a particular family of posets which we call Wixárica posets. We will then describe how topological information of these posets is transferred to the algebraic level.
The remainder is organized as follows. In Section 2 we describe posets and introduce Wixárika posets. In Section 3 we introduce an algebra over the operad of Wixárika posets: the order series. Section 4 formalizes the notion of operads and their algebras. Section 5 provides a worked-out example in enumerative combinatorics. Section 6 provides an example of the use of topological information that is transferred among algebras over an operad. We then suggest future reading in Section 7.
2. Wixárica Posets
In this section, we will define our main object of study, which is the class of Wixárica posets. We are interested in some algebraic properties of these posets. In particular, we will define the operations of concatenation and disjoint union of posets. Moreover, we will introduce a new operation, called the ’itsari operation, to obtain Wixárica posets through the consecutive application of these operations. We will look at the Hasse diagrams of Wixárica posets in order to have graphical representations.
When dealing with posets, we have two definitions: the strict and the non-strict orders. Most of the time, we are working with the second. We shall give the two for completeness.
Definition 1.
A partial order $\leq$ on a set $P$ is a reflexive, antisymmetric and transitive relation on its elements. That is, for $a,b,c\in P$:
  1. 1)
    Reflexivity: $a\leq a$, i.e., every element is related to itself.
  2. 2)
    Antisymmetry: if $a\leq b$ and $b\leq a$, then $a=b$, i.e., no two distinct elements precede each other.
  3. 3)
    Transitivity: if $a\leq b$ and $b\leq c$, then $a\leq c$.
Definition 2.
A strict partial order << on a set $P$ is an irreflexive, asymmetric and transitive relation on its elements. That is, for $a,b,c\in P$:
  1. 1)
    Irreflexivity: no element $a$ is related to itself.
  2. 2)
    Asymmetry: if $a<b$, then it is not possible that $b<a$.
  3. 3)
    Transitivity: if $a<b$ and $b<c$, then $a<c$.
Definition 3.
A set $P$ together with a (strict) partial order $(P,\leq)$ is called a (strict) partially ordered set or simply a (strict) poset. We use the notation for the set and the poset with underlying set $P$. To avoid confusion, we use $\leq_{P}$ to refer to implicit order.
Example 4.
Examples of posets. Ordered sets, such as the real numbers; the set of subsets of a given set, its power set, ordered by inclusion; the set of natural numbers equipped with the divisibility relation; if we consider a finite set of programs with the order relations $A<B$ whenever the output of $A$ is an input of $B$, then a set of programs forms a poset where the order gives restrictions to select at what turn to run each program.
In this work we assume every poset is finite.
A Hasse diagram is a graph used to represent a finite partially ordered set, in the form of a drawing of its transitive reduction. We will draw Hasse diagrams when we need to visualize certain posets. Concretely, a partially ordered set $P$ with order $\leq$ is represented by a graph whose vertices are the elements of $P$, and we have an edge from one vertex $r$ to another vertex $s$ whenever $r\neq s$ and $r<s$ and there is no $t$ with $r<t<s$. Typically, in Hasse diagrams, the greater element is positioned higher than the lesser.
Example 5.
An example is given in Figure 1, it shows the poset given by the divisors of 60.
Figure 1. Hasse diagram of the poset of divisors of 60 (Image created by Ed_g2s, https://commons.wikimedia.org/wiki/User:Ed_g2s).
We use a particular class of posets, series-parallel posets. These posets are defined constructively, starting from a set of points and iterated applications of operations that we now introduce.
Definition 6.
We define the direct sum of two posets $(P_{1},\leq_{P_{1}})$ and $(P_{2},\leq_{P_{2}})$ to be the poset $P=P_{1}*P_{2}$ whose underlying set is the union of $P_{1}\sqcup P_{2}$ and $a\leq_{P}b$ if and only if
  1. 1)
    $a,b\in P_{1}$ and $a\leq_{P_{1}}b$, or
  2. 2)
    $a,b\in P_{2}$ and $a\leq_{P_{2}}b$, or
  3. 3)
    $a\in P_{1}$ and $b\in P_{2}$.
Definition 7.
The disjoint union poset of $(P_{1},\leq_{P_{1}})$ and $(P_{2},\leq_{P_{2}})$ is the poset whose underlying set is $P_{1}\sqcup P_{2}$ and there are no more relations than those given by $\leq_{P_{1}}$ and $\leq_{P_{2}}$.
In the context of series-parallel posets, the direct sum is called series composition while parallel composition refers to the disjoint union.
Definition 8.
The class of series-parallel orders is defined inductively as follows:
  1. 1)
    A single element is series-parallel.
  2. 2)
    If $(P_{1},\leq_{P_{1}})$ and $(P_{2},\leq_{P_{2}})$ are series parallel with $P_{1}\cap P_{2}=\emptyset$, then their series composition is series-parallel.
  3. 3)
    If $(P_{1},\leq_{P_{1}})$ and $(P_{2},\leq_{P_{2}})$ are series-parallel with $P_{1}\cap P_{2}=\emptyset$, then their parallel composition is series-parallel.
Example 9.
An example is illustrated in Figure 2.
Figure 2. Hasse diagram of a series-parallel poset (Image created by David Eppstein, https://commons.wikimedia.org/wiki/User:David_Eppstein/Gallery).
Next, we introduce a new operation on single posets. Consecutive applications of this operation and disjoint union will produce the Wixárica posets. Note that Wixárica posets are a special case of series-parallel posets.
Definition 10.
The ’itsari operation $D$ is defined as
$D(P)=\langle 1\rangle*\big(\langle 1\rangle\sqcup P\big)*\langle 1\rangle.$
’Itsari is a word from Wixárica language whose meaning is to weave. Geometrically, we are adding a new minimum $x_{0}$ and a new maximum $x_{1}$ to $X$, then we attach a handle on these points, that is, we add a third element $y$ with $x_{0}<y<x_{1}$.
Definition 11.
A Wixárika poset is a finite poset obtained from the poset with one element by finitely many applications of the operations $D$ and *.
Example 12.
An example of a Wixárica poset is shown by the Hasse diagram in Figure 3. Note that in contrast to the previous examples, we direct Hasse diagrams of Wixárica posets from left to right.
Figure 3. Hasse diagram of a Wixárika poset.
The Hasse diagrams of these posets resemble Wixárika collars, which are intricate colorful necklaces crafted by the Wixárika community in North America.
3. Order Series
In this section, we review the theory of power series and describe a set of power series associated to posets, as well as operations between them. The idea is that the operations introduced in this section are related to the poset operations defined in the previous section. Most of our examples of posets and ordered sets are given using a particular type of very simple poset, called a chain.
Definition 13.
Define a $n$-chain to be the total ordered set $\langle n\rangle=\{1<2<\cdots<n\}$ and let $P$ be a finite partially ordered set. An order-preserving map $\sigma:P\to\langle n\rangle$ is a map such that if $r\leq s$ in $P$, then $\sigma(r)\leq\sigma(s)$. Define $\Omega^{\circ}(P,n)$ as the number of order-preserving maps from $P$ to $\langle n\rangle$. A strictly order-preserving map $\tau:P\to\langle n\rangle$ is a map such that if $r<s$ then $\tau(r)<\tau(s)$. Define $\Omega(P,n)$ to be the number of strictly order-preserving maps from $P$ to $\langle n\rangle$.
Definition 14.
[…]
Example 15.
[…]
Definition 16.
For a poset $P$ we define a formal power series, called the order series of $P$, by
$\mathfrak{Z}\left(P\right)=\mathfrak{Z}\left(P,x\right)=\sum_{n=1}^{\infty}\Omega^{\circ}(P,n)x^{n}.$
Example 17.
For instance, we have
(1) $\mathfrak{Z}\left(m\right):=\mathfrak{Z}\left(\langle m\rangle\right)=\sum_{n=m}^{\infty}{n\choose m}x^{n}=\frac{x^{m}}{(1-x)^{m+1}}.$
For a proof of the second equality see [WIL06, Equation (1.5.5)], [GOU72, Equation (1.3)] or [PET15, Equation (1.3)].
We will use some standard operations on power series with order series. We include them for completeness.
Definition 18.
If $\sum_{i=0}^{\infty}a_{i}x^{i}$ and $\sum_{i=0}^{\infty}b_{i}x^{i}$ are two power series, the Cauchy product is defined as
$\left(\sum_{i=0}^{\infty}a_{i}x^{i}\right)\left(\sum_{i=0}^{\infty}b_{i}x^{i}\right)=\sum_{n=0}^{\infty}c_{n}x^{n}\quad\text{where}\quad c_{n}=\sum_{k==0}^{n}a_{k}b_{n-k}.$
Definition 19.
If $x$ is the common variable for the order series $\mathfrak{Z}\left(Q\right)$ and $\mathfrak{Z}\left(P\right)$, we define the direct product using the Cauchy product as follows:
(2) $\mathfrak{Z}\left(Q\right)*\mathfrak{Z}\left(P\right)=\mathfrak{Z}\left(Q\right)(1-x)\mathfrak{Z}\left(P\right).$
Definition 20.
We denote the Hadamard product of power series by
$\sum_{n=1}^{\infty}a_{n}x^{n}\odot\sum_{n=1}^{\infty}b_{n}x^{n}=\sum_{n=1}^{\infty}a_{n}b_{n}x^{n}.$
In the case of order series we will call the Hadamard product disjoint union and use the symbol $\sqcup$, that is,
$\mathfrak{Z}\left(P\right)\sqcup\mathfrak{Z}\left(Q\right)=\mathfrak{Z}\left(P\right)\odot\mathfrak{Z}\left(Q\right).$
Definition 21.
We define the ’itsari operation for power series as
$D(\mathfrak{Z}\left(P\right))=\mathfrak{Z}\left(D(P)\right).$
For the case of chains $\langle n\rangle$ we denote $D(\langle n\rangle)=D(n)$.
Note that we are abusing notation: We are using the same symbol for operations between posets and order series. The meaning, however, will be clear from the context.
The following two results are well-known (see [FS86], [WAG92], [BRA06], [BHK18],[ABD23]).
Proposition 22.
If $P$ and $Q$ are two posets, then
$\mathfrak{Z}\left(P*Q\right)=\mathfrak{Z}\left(P\right)(1-x)\mathfrak{Z}\left(Q\right).$
Proposition 23.
For two chains we have
$\mathfrak{Z}\left({\langle k\rangle}\sqcup{\langle m\rangle}\right)=\sum_{n=0}^{k}{m+n\choose k}{k\choose n}\mathfrak{Z}\left(m+n\right).$
Corollary 24.
The strict order series of any series-parallel poset is a linear combination of order series of chains.
The following result will be very useful for one of our main objectives, which is to describe the order series of any series-parallel poset.
Proposition 25.
Unraveling the definitions, we can show linearity of $D$:
$D(a\mathfrak{Z}\left(P\right)+b\mathfrak{Z}\left(Q\right))=aD(\mathfrak{Z}\left(P\right))+bD(\mathfrak{Z}\left(Q\right)).$
Example 26.
If the input of the ’itsari operation $D$ is an $n$-chain, we compute
$\mathfrak{Z}\left(D({\langle n\rangle})\right)$ $=\mathfrak{Z}\left(1\right)*\mathfrak{Z}\left(\langle 1\rangle\sqcup\langle n\rangle\right)*\mathfrak{Z}\left(1\right)$
$=\frac{x}{(1-x)^{2}}*(n\frac{x^{n}}{(1-x)^{n+1}}+(n+1)\frac{x^{n+1}}{(1-x)^{n+2}})*\frac{x}{(1-x)^{2}}$
$=n\frac{x^{n+2}}{(1-x)^{n+3}}+(n+1)\frac{x^{n+3}}{(1-x)^{n+4}}$
$=n\mathfrak{Z}\left(n+2\right)+(n+1)\mathfrak{Z}\left(n+3\right).$
4. Algebras over the operad of Wixárika posets
In this section, we introduce the language of operads. This allows us to compare sets in which the same operations are defined, similarly to homomorphisms of groups.
Definition 27.
An operad $(\mathcal{O},\circ)$ consists of a collection $\mathcal{O}=\{O(n)\}_{n\geq 1}$ of sets and an associative composition $\circ:O(n)\times O(k_{1})\times\cdots\times O(k_{n})\to O(\sum_{i=1}^{n}k_{i})$. There are unital elements $id\in O(1)$, and there is an action of $S_{n}$ that is compatible with the composition operation.
In the composition, an element in $O(n)$ takes the role of a map with $n$ inputs, and an element in $O(k_{1})\times\cdots\times O(k_{n})$ represents $n$ maps that are composed with the former. The result is a map with $\sum_{i=1}^{n}k_{i}$ inputs .
Definition 28.
Given an operad $\mathcal{O}$ and $f\in O(n)$, for every $i\leq n$ and $g\in O(k_{i})$ the partial composition $\circ_{i}$ is defined by $f\circ_{i}g=f\circ(id,\cdots,id,g,id,\cdots,id)$ where $g$ is in the position of the $i$-th input of $f$.
A more detailed definition of operad can be found in  [STA04, CHA07, LV12, VAL14, NAO20, BRA21, HM22].
Example 29.
The most fundamental example of an operad is the endomorphism operad $\mathrm{End}_{A}$: Let $A$ be a set and $\mathrm{Map}(A^{n},A)$} the set of functions from the cartesian product $A^{n}$ to $A$. Then $\mathrm{End}_{A}$ has as $n$-ary elements $\mathrm{End}_{A}(n)={\mathrm{Map}(A^{n},A)}$.
Example 30.
Consider the tree grafting operad: Let $T(n)$ be the set of trees with $1$ root and $n$ leaves labeled (arbitrarily) $1$ through $n$. The collection $\{T(n)\}$ forms an operad by grafting the root of $g$ to the leaf of $f$ labeled $i$, as in Figure 4.
Figure 4. Operations on the operad tree (Image created by Wiki_cies, https://commons.wikimedia.org/w/index.php?title=User:Wiki_cies&action=edit&redlink=1).
Definition 31.
By an algebra $A$ over an operad $\mathcal{O}$ we mean an operadic morphism $\mathcal{O}\mapsto\mathrm{End}_{A}$.
An algebra $A$ over an operad $\mathcal{O}$ is then a set in which the abstract elements of $O(n)$ are realized as functions $A^{n}\mapsto A$. Heuristically, we can think of elements $\sigma\in O(n)$ of an operad as $n$-ary trees. Then a composition $\sigma\circ_{i}\tau$ corresponds to the grafting of the trees were we identify the leaf of $\sigma$ with the root of $\tau$. The evaluation $\sigma(a_{1},\cdots,a_{n})$ where $a_{1},\cdots,a_{n}$ are elements of $A$ then corresponds to placing drops of water on the leaves of the tree. The drops fall along the tree. Whenever a drop reaches a branching point (labeled with an operation), we evaluate the corresponding operation and replace the drop by the output. The process continues until we reach the root.
Definition 32.
The Wixárica operad is the operad $W$ generated by points under the binary operation $\ast$ and the unary operation $D$.
Example 33.
The elements of the operad $W$ can be described by binary rooted trees with bird nests: There are vertices with two incoming edges and one outgoing edge labeled by *; and the bird nest stand for vertices with one incoming and one outgoing edge, which are labeled by $D$. See Figure 5.
Figure 5. Left: The identity operation, the operation $D$ and the operation $\ast$. Right: A tree representing the element in Figure 3.
The composition of posets then allows us to see Wixárika posets as algebras over the operad $W$.
Example 34.
Consider Figure 6. On the left side we have at the top three copies of the posets $\langle 1\rangle$. Below is a sequence of transformations using $D$ and *. At the bottom we have the final result, a Wixárika poset. On the right side we have the corresponding tree of operations in which the leaves should be thought of as decorated by the poset $\langle 1\rangle$. The vertices with two incoming edges and one outgoing edge are labeled with * and the vertices with one incoming edge and one outgoing edge (the bird nest) are labeled with $D$.
Figure 6. To the left bottom we have a Wixárika poset, as an element of the algebra. To the right we have the tree of operations, an element of the operad of Wixárika posets. When this element is evaluated on six points it produces the Wixárika poset of the left. As the drops of water run down the tree, the poset on the left is being assembled.
Definition 35.
Given two algebras $A,B$ over an operad $\mathcal{O}$, an operadic morphism $\phi:A\to B$ is a map that makes the following diagram commutative:
[…]
Given $P$ a Wixárika poset, define
$D(\mathfrak{Z}\left(P\right)):=\mathfrak{Z}\left(D(P)\right).$
Proposition 36.
The function $\mathfrak{Z}\left(\cdot\right)$ defines an operadic homomorphism between the $W$-algebra of Wixárika posets and the $W$-algebra of order series of Wixárika posets.
Proof.
Put together Definition 19 and Proposition 22 to get:
$\mathfrak{Z}\left(P\ast Q\right)=\mathfrak{Z}\left(P\right)(1-x)\mathfrak{Z}\left(Q\right)=\mathfrak{Z}\left(P\right)\ast\mathfrak{Z}\left(Q\right).$
By Proposition 23 and its Corollary 24 we have:
$\mathfrak{Z}\left(P\sqcup Q\right)=\mathfrak{Z}\left(P\right)\sqcup\mathfrak{Z}\left(Q\right).$
We obtain the explicitly formula
$D(\mathfrak{Z}\left(P\right))=\mathfrak{Z}\left(1\right)\ast(\mathfrak{Z}\left(P\right)\sqcup\mathfrak{Z}\left(1\right))\ast\mathfrak{Z}\left(1\right).$
5. Evaluation on order series
In this section, we give an explicit example of how to compute the order series of a Wixárica poset using the algebra homomorphims between both $W$-algebras defined in the previous section. To compute the order series, we follow [FS86], but work on the level of power series instead of order polynomials.
Consider the homomorphism between the algebra over the operad $W$ generated by $\langle 1\rangle$ and the algebra generated by $\mathfrak{Z}\left(1\right)$. In Proposition 36 we constructed $\ast,\sqcup$ such that
$\mathfrak{Z}\left(P\right)*\mathfrak{Z}\left(Q\right)=\mathfrak{Z}\left(P*Q\right),\quad D(\mathfrak{Z}\left(P\right))=\mathfrak{Z}\left(D(P)\right).$
If we have an element of the operad, which we describe as a binary tree with bird nests, then we obtain a poset by placing points $\langle 1\rangle$ on the leaves of the tree. Similarly, we obtain a power series by placing $\mathfrak{Z}\left(1\right)$ on the leaves of the tree, that is, $\mathfrak{Z}\left(\right)$ substitutes any input leaf $\langle 1\rangle$ by $\mathfrak{Z}\left(1\right)$.
Example 37.
Consider the poset in Figure 6. The resulting poset is given by
$P=D(D(D(D(1)*\langle 1\rangle*D(1)))*\langle 3\rangle).$
We shall give a procedure to calculate $\mathfrak{Z}\left(P\right)$. The first step of our algorithm is replacement of the leaves $\langle 1\rangle$ by their order series $\mathfrak{Z}\left(1\right)=\frac{x}{(1-x)^{2}}$. In the tree, going from top to bottom, the algebra homomorphism allows us to make the following replacements: a concatenation PQP*Q by $\mathfrak{Z}\left(P\right)*\mathfrak{Z}\left(Q\right)$ and $D(n)$ by $\mathfrak{Z}\left(D(n)\right)$ in $P$. Using Proposition 22 and Example 26, as well as linearity of $D$ and bilinearity of $\ast$, one evaluates the vertex operations until we reach the root to obtain
$D(D(D(D(1)*\mathfrak{Z}\left(1\right)*D(1)))*\mathfrak{Z}\left(3\right))$
$=D(D(D((\mathfrak{Z}\left(3\right)+2\mathfrak{Z}\left(4\right))*\mathfrak{Z}\left(1\right)*(\mathfrak{Z}\left(3\right)+2\mathfrak{Z}\left(4\right))))*\mathfrak{Z}\left(3\right))$
$=D(D(D(\mathfrak{Z}\left(7\right)+4\mathfrak{Z}\left(8\right)+4\mathfrak{Z}\left(9\right)))*\mathfrak{Z}\left(3\right))$
$=D(D(D(\mathfrak{Z}\left(7\right)))*\mathfrak{Z}\left(3\right))+4D(D(D(\mathfrak{Z}\left(8\right)))*\mathfrak{Z}\left(3\right))$
$+4D(D(D(\mathfrak{Z}\left(9\right)))*\mathfrak{Z}\left(3\right))$
This is an illustrative algorithm, remember that we are abusing notation: $\ast$ and $D$ for concatenation and ’itsari operations must be applied in their respective algebra. This is justified by the algebra homomorphism. Following the same procedure we continue
$882\frac{x^{16}}{(1-x)^{17}}+7995\frac{x^{17}}{(1-x)^{18}}+27232\frac{x^{18}}{(1-x)^{19}}+$
$43792\frac{x^{19}}{(1-x)^{20}}+33552\frac{x^{20}}{(1-x)^{21}}+9880\frac{x^{21}}{(1-x)^{22}}.$
As a second step, to compute the order polynomial $\Omega^{\circ}(m)$, we use the expression $\frac{x^{n}}{(1-x)^{n+1}}=\sum_{m\geq n}{m\choose n}x^{m}$ to extract the term of degree $x^{m}$ from the order series. Because our work focuses on studying operations that preserve the number of labeling maps of posets, we know that the coefficient of $x^{m}$ is the number of labelings of the poset using numbers from $1$ to $m$.
From the above we thus find that for any $m$ the number of labeling maps of $P$ is given by
$882{m\choose 16}+7995{m\choose 17}+27232{m\choose 18}+43792{m\choose 19}+33552{m\choose 20}+9880{m\choose 21}$.
6. Representability
The Wixárika posets were selected due to their topological properties. We now use the topological information to answer the inverse problem: Given a function $f(x)$, does there exist a poset $P$ whose order series coincides with $f(x)$? In this section, we present an algorithm to answer this question.
Definition 38.
We say a function $f(x)$ is represented if there is a Wixárika poset $P$ such that $f(x)=\mathfrak{Z}\left(P\right)$.
Example 39.
Two posets are called Doppelgänger posets if they share the same order series, for example $\langle 1\rangle*(\langle 1\rangle\sqcup\langle 1\rangle\sqcup\langle 1\rangle)$ and $\langle 2\rangle\sqcup\langle 2\rangle$ have the same order series: $\mathfrak{Z}\left(2\right)+6\mathfrak{Z}\left(3\right)+6\mathfrak{Z}\left(4\right)$ [HPP+18].
In order to decide whether a function can be represented, we determine topological invariants of a poset $P$ from its associated power series $\mathfrak{Z}\left(P\right)$, such as the Betti number of the Hasse diagram of $P$ (i.e., the number of independent cycles in the underlying undirected graph). With this information, we describe an algorithm that determines whether $f(x)$ can be represented. If this is the case, then the algorithm explicitly constructs all possible Wixárika posets $P$ such that $f(x)=\mathfrak{Z}\left(P\right)$.
The following theorem states that the coefficients of the order series associated to a Wixárika poset are topological invariants of its Hasse diagram.
Theorem 40.
Let $P$ be a Wixárika poset. Consider its factorization, that is, the associated word $w$ in the characters * and $D$. Let $f(x)=w(\mathfrak{Z}\left(1\right))$ be the power series obtained by replacing every leaf of $w$ by $\mathfrak{Z}\left(1\right)$ and proceeding as in Example 37.
  1. (1)
    The series $f(x)$ admits a unique description
    $f(x)=d_{i}\mathfrak{Z}\left(i\right)+d_{i+1}\mathfrak{Z}\left(i+1\right)+\cdots+d_{k}\mathfrak{Z}\left(k\right),$
    where the coefficients $d_{j}$ are non-negative integers.
  2. (2)
    The number $i$ is the number of points in a maximal chain in $P$.
  3. (3)
    The number $k$ is the number of points in $P$.
  4. (4)
    The difference d=kid=k-i is the first Betti number of the Hasse diagram of $P$, equal to the number of times $D$ occurs in $w$.
  5. (5)
    The difference m=i2d1m=i-2d-1 is the number of times * occurs in $w$.
  6. (6)
    The term $m+1$ is the number of leaves in the tree $w$.
  7. (7)
    $\sum_{u=1}^{k}(-1)^{k-u}d_{u}=1$.
Sketch of the proof.
One shows that every Wixárika poset has a unique minimum and a unique maximum. Then all statements follow by describing the possible values obtained under the operations $D$ and *. The explicit description of $D$ and * leads to the positivity of the coefficients.
Consider, for example, statement $(4)$: We know that by applying $D$ we increase the first Betti number of the Hasse diagram of a poset by 1. The first Betti number of a concatenation is the sum of the first Betti numbers of each element, because Wixárika posets have a unique maximum and a unique minimum. Since Wixárika posets are generated by $\langle 1\rangle$, and the Betti number of $\langle 1\rangle$ is zero, it follows that the only way to increase the Betti number of a poset is by the operation $D$ or concatenating the poset with another poset with non-zero Betti number.
For details, see [ABD23, Proposition 3.1]. ∎
Example 41.
The series
$882\frac{x^{16}}{(1-x)^{17}}+7995\frac{x^{17}}{(1-x)^{18}}+27232\frac{x^{18}}{(1-x)^{19}}+$
$43792\frac{x^{19}}{(1-x)^{20}}+33552\frac{x^{20}}{(1-x)^{21}}+9880\frac{x^{21}}{(1-x)^{22}},$
corresponds to the poset depicted on the left bottom of Figure 6. We visually verify that the poset has $i=16$ points in any chain of maximal size, consists of $k=21$ elements in total, the Betti number of its Hasse diagram equals ki=5k-i=5, is generated by m=162(5)1=5m=16-2(5)-1=5 uses of *, and its tree on the right has $6$ leaves.
The paper [BHK18] calls the elements $\{\mathfrak{Z}\left(i\right)\}_{i\in\mathbb{N}}$ the chain basis. The coefficients of order series in the chain basis satify the following positivity property:
Theorem 42.
Consider $\mathfrak{Z}\left(P\right)=\sum_{i=j_{0}}^{|P|}d_{i}\mathfrak{Z}\left(i\right),d_{i}\geq 0$. The order polytope of $P$ is the union of $d_{|P|}$ copies of the $|P|$-simplex. The remaining coefficients $d_{i}$ of $\mathfrak{Z}\left(P\right)=\sum d_{i}\mathfrak{Z}\left(i\right)$ count the number of $i$-simplices that occur, in the canonical triangulation of the order polytope, as intersection of the $d_{|P|}$ copies of those $|P|$-simplices.
Proof.
See [ABD23, Proposition 3.6]. ∎
The coefficients of power series associated to more general posets, for instance series-parallel posets (which are generated by concatenation and disjoint union), are not topological invariants. For example, the Hasse diagram of the poset $\langle 1\rangle*(\langle 1\rangle\sqcup\langle 1\rangle)*\langle 1\rangle$ has first Betti number $1$ while the Hasse diagram of the poset $\langle 2\rangle*(\langle 1\rangle\sqcup\langle 1\rangle)$ has first Betti number 0. Their order series, however, are the same.
In the other direction, we have the following statement.
Theorem 43.
Given a finite sum $f(x)=\sum d_{j}\mathfrak{Z}\left(j\right)$ where $d_{j}$ are non-negative integers, there is an algorithm to determine if $f(x)=\mathfrak{Z}\left(P\right)$ for some Wixárika poset $P$. In the positive case the algorithm returns all possible posets representing $f(x)$.
Sketch of the proof.
We first verify that the function $f(x)$ satisfies the requirements of Theorem 42. Then we restrict the search space using Theorem 40. Using colored operads, it is possible to reduce the search to a set $C$ of representatives of equivalence classes of trees up to associativity. Now for every representative $P\in C$ we first evaluate the first and last term of $\mathfrak{Z}\left(P\right)=b_{i}\mathfrak{Z}\left(i\right)+\cdots+b_{k}\mathfrak{Z}\left(k\right)$. And this evaluation can be optimized as well since we know what are the final values. If we find a poset $Q$ with $\mathfrak{Z}\left(Q\right)=r_{i}\mathfrak{Z}\left(i\right)+\cdots+r_{k}\mathfrak{Z}\left(k\right)$ and $b_{i}=r_{i},b_{k}=r_{k}$ then we compute all terms ${\{r_{j}\}}_{i<j<k}$ and compare with $f(x)$ until we find a solution of we finish the candidate posets. The evaluation can be done with the tree of operations as described in [FS86, Section 4], but working with power series instead of polynomials. If we find a poset that produces $f(x)$, we use the results of [BHK18, Sections 4.2 and 5 ] to find the remaining Doppelgängers, that is, posets that have the same power series $f(x)$. After this step, we stop the algorithm. For details, see [ABD23, Proposition 3.2]. ∎
As a filter, one can also verify a condition coming from Ehrhart theory  [15].
Remark 44.
Theorem 43 requires a function already written as a finite linear combination of terms $\{\mathfrak{Z}\left(n\right)\}_{n\in\mathbb{N}}$. The reason is that a random power series may not be expressable as a finite linear combination of the $\{\mathfrak{Z}\left(n\right)\}_{n\in\mathbb{N}}$. For example, consider the function $\sum_{n=1}^{\infty}\sum_{j=1}^{n}{n\choose j}x^{n}$.
Example 45.
Consider the function $8\mathfrak{Z}\left(13\right)+25\mathfrak{Z}\left(14\right)+18\mathfrak{Z}\left(15\right)$. The algorithm of Theorem 43 efficiently tests posets until it finds a candidate. Suppose the algorithm finds $Q=D(\mathfrak{Z}\left(8\right))*D(\mathfrak{Z}\left(1\right))$. Then with [BHK18, Section 5.3] we obtain $Q_{2}=D(D(\mathfrak{Z}\left(2\right)))*\mathfrak{Z}\left(7\right)$ as a Doppelgänger of $Q$. Up to commutativity of *, these are all the Doppelgänger Wixárika posets.
7. Farewell
In this note, we aimed to show how operads can relate objects of different nature such as posets and power series. We also gave an example in which the topological information of the operad can be used to improve an algorithm.
What other operads can interested readers study? There is an encyclopedia of operads, a project in progress by Frédéric Chapoton: Operadia https://operadia.pythonanywhere.com/table, and the Encyclopedia of types of algebras [ZIN10]. If one aims to study a new operad, the key point is to have a notion of composition, for example, cylinders can be composed vertically. A variation of the operad of cylinders encodes all the operations one expects in calculus, as described in [DTT08, Theorem 2].
The following list is not exhaustive, but a sample to illustrate papers with applications of operad theory. After all, in February 2026 the website zbMATH returns 264 papers with the word ”operad” in the title. Operad theory started with the work of J. M. Boardman, R. M. Vogt [BV68], and P. May [MAY72], in the field of algebraic topology. M. Kontsevich [KON99], D. Tamarkin [TAM03], and T. Willwacher [ŠW11] popularized them in mathematical physics. F., Chapoton [CHA07] described the relationship between operads and combinatorial species, S. Giraudo [GIR17] studies combinatorial problems as algebras of some operads. F. Fauvet, L. Foissy, and D. Manchon [FFM18] studied the operadic structures on finite posets. J. Baez, J. Foley, J. Moeller, and B. S. Pollard studied network models via operads in [BFM+20].
The reader may be surprised by the diversity of fields in which operad theory is used. This is however quite natural, since operads are a categorical concept, and the aim of category theory is to study mathematics from a universal point of view.
8. Acknowledgments
An old version of these notes was used to give a minicourse in Govt. College University Lahore in Pakistan 2022, and at the BIRS-CMO workshop “6th meeting of the Mexican Mathematicians in the World” 2024. The work of E. Dolores Cuenca was supported by the Korea National Research Foundation (NRF) grant funded by the Korean government (MSIT) (RS-2025-00517727).